# Beyond Complete Shapes: A Quantitative Evaluation of 3D Shape Matching Algorithms

### v1

@@ Abstract
Finding correspondences between 3D shapes is an important and long-standing problem in computer vision, graphics and beyond. While approaches based on machine learning dominate modern 3D shape matching, almost all existing (learning-based) methods require that at least one of the involved shapes is complete. In contrast, the most challenging and arguably most practically relevant setting of matching partially observed shapes, is currently underexplored. One important factor is that existing datasets contain only a small number of shapes (typically below 100), which are unable to serve data-hungry machine learning approaches, particularly in the unsupervised regime. In addition, the type of partiality present in existing datasets is often artificial and far from realistic. To address these limitations and to encourage research on these relevant settings, we provide a generic and flexible framework for the procedural generation of challenging partial shape matching scenarios. Our framework allows for a virtually infinite generation of partial shape matching instances from a finite set of shapes with complete geometry. Further, we manually create cross-dataset correspondences between seven existing (complete geometry) shape matching datasets, leading to a total of 2543 shapes. Based on this, we propose several challenging partial benchmark settings, for which we evaluate respective state-of-the-art methods as baselines. Our entire benchmark can be viewed on: https://becos-authors.github.io/BeCoS/.
Figure 1: Our framework generates correspondences (depicted via colour transfer from the centaur on stage) between thousands of shapes. The gray parts encode the missing parts of partial shapes. With our framework we can generate virtually unlimited partial shape matching data across diverse four-legged animals and humanoid categories.
footnotetext: : These authors contributed equally to this work.
@@ 1 Introduction
Finding correspondences between deformable shapes is a fundamental problem in computer vision and graphics, with various applications such as texture and deformation transfer [19, 42], or statistical shape analysis [30, 10, 27]. The problem has been widely studied in the literature  [42, 25, 1, 54, 15], and can vary significantly in complexity. For example, current methods still struggle when estimating severe non-isometric deformations for shapes of different categories [22, 44, 25, 7, 16], or when they are only partially observed [48, 24]. In real-world scenarios, it is particularly common to encounter shapes with missing parts due to occlusions, limited sensor coverage, or acquisition errors during 3D scanning. Furthermore, a semantic or geometric similarity might exist only between parts of the shapes but not between their whole geometry.
@@
The setting in which at least part of one shape does not exist on the other is called partial shape matching [46, 3, 48, 23, 24]. Despite its practical relevance, partial shape matching has received limited attention in machine learning, partly because of increased difficulty, but largely due to the lack of suitable and realistic large-scale datasets, a requirement for data-hungry learning algorithms. The full-to-full matching setting is not only simpler but also benefits from wider data availability, which has supported the development of data-driven methods that already display convincing performance even without supervision [38, 20, 15]. In contrast, existing datasets for partial shape correspondence often suffer from a small number of shapes, limited realism of generated partiality (e.g., cuts or holes) [17], and a lack of diversity in shape categories [11, 3]. This limited data availability hampers the progress of data-driven methods for partial shape matching; recent approaches have thus required pre-training with large full-geometry datasets [15], or rely on ground-truth correspondences [3] which can be challenging to obtain in practice.
@@
To address the absence of large partial datasets and foster research on partial shape correspondence, we introduce a procedural data generation framework to generate realistic and diverse partial shape pairs. Our framework is designed to unify existing full geometry correspondence datasets to generate virtually infinite partial shape pairs with known correspondences. By simulating realistic partiality patterns using ray casting, we bridge the gap between existing synthetic datasets and real-world partial shape correspondence problems. Furthermore, we propose BeCoS, a carefully curated benchmark based on existing datasets comprising diverse shape categories, and containing more realistic partial instances to assess the performance of existing and foster the development of novel methods. Through extensive experiments, we demonstrate the limitations of current state-of-the-art methods for partial shape correspondence and highlight the need for further research in this area.
@@
  • A flexible and scalable framework that enables researchers to generate virtually unlimited training examples for partial shape correspondence tasks based on existing (full or partial geometry) data. The framework supports the selection and propagation of cross-category matchings and provides a versatile solution for creating custom training data tailored to specific application requirements.
  • A challenging partial non-rigid shape dataset constituting a carefully curated instantiation of our framework. This dataset serves as a standardized benchmark for shape matching methods, promoting the development of more robust and accurate algorithms.
  • An analysis of the performance of current state-of-the-art methods on the new benchmark, providing valuable insights into their performance, limitations, and key challenges. Results suggest that partial shape correspondence is far from being solved.
@@ 2.1 Datasets
In the following, we provide an overview of existing datasets for partial-to-full (P2F), and partial-to-partial (P2P) shape matching datasets. Table 1 summarises existing partial shape matching datasets.
Dataset # Unique Shapes Shape Type
P2F SHREC’16 [13, 17] 76 [Uncaptioned image]/[Uncaptioned image]
PFAUST [11, 10] 10 [Uncaptioned image]
FARM [34] 5 [Uncaptioned image]
PFARM [3, 34] 28 [Uncaptioned image]
SHREC’20 [22]* 14 [Uncaptioned image]
P2P CP2P [3, 17] 76 [Uncaptioned image]
PARTIALSMAL [24, 56]* 43 [Uncaptioned image]
Ours* 2543 [Uncaptioned image]/[Uncaptioned image]
Table 1: An Overview of Partial 3D Shape Datasets, categorized as partial-to-full (P2F) and partial-to-partial (P2P), including the number of available shapes shape type (Humanoid [Uncaptioned image]or Four-legged [Uncaptioned image]), and deformation class (Near-Isometric and Non-Isometric (*)). Due to limited correspondence pairs, FARM and PFARM are typically only used for evaluation.
Full-to-Full Datasets.
@@
Full-to-full shape correspondence datasets, such as FAUST [10] and SCAPE [2], have been essential in methodological progress on shape correspondence. They encompass humans in various poses, initially emphasising point-wise correspondence recovery under isometric deformations. SHREC’19 [40] was introduced to evaluate the generalisation capabilities of methods to different mesh connectivity on human scans. Datasets like TOSCA [13], SMAL [56], SHREC’20 [22], and DeformingThings4D [39, 36] have shifted the emphasis towards non-isometric correspondence between various animal categories and humanoid figurines.
Partial-to-Full Datasets.
While full-to-full shape matching has received the most attention, several partial shape matching datasets have also been proposed, in which a partial shape is deformably matched to a shape template (with complete geometry). SHREC’16 [17], based on the TOSCA [13] dataset, was the first benchmark for this problem setting. PFAUST [11] and PFARM [3] were later introduced using a similar approach based on the FAUST [10] and FARM [34] datasets, respectively. SHRE C’20 [22] also contains a subset of full-to-partial correspondence pairs.
Partial-to-Partial Datasets.
Partial-to-partial shape matching involves matching two partial shapes with unknown overlap, making it considerably more challenging than partial-to-full matching. Attaiki et al. proposed the first benchmark for this setting with CP2P [3] based on the SHREC’16 [17] CUTS benchmark. Recently, Ehm et al. introduced PARTIALSMAL [24], extending the SMAL dataset to include partial-to-partial correspondence pairs.
2.2 Methods for Shape Correspondence
We proceed by discussing a selection of shape matching methods and refer to recent surveys  [50, 18] for a more thorough treatment. While point-cloud matching methods also exist [35, 6, 32, 55], we focus on mesh-based methods, as our dataset is designed for them. Methods addressing shape correspondence can be categorized into axiomatic and learning-based. Among the axiomatic approaches are discrete optimization methods [8] and spectral methods utilising the powerful framework of functional maps [42, 46, 41]. However, the performance of the latter class degrades as assumptions of isometry are violated, such as in partial correspondence settings, which we emphasize in this work.
Full-to-Full Shape Matching.
Spectral methods based on the eigenfunctions of the Laplace-Beltrami Operator (LBO) have been very successful in full-to-full matching settings [42, 38, 21, 7, 15]. Functional maps simplify the challenging quadratic assignment problem in shape matching to a linear change of basis in the LBO eigenfunctions, yielding a compact solution [42]. In contrast to spectral approaches, discrete optimization formulations have also been introduced for full-to-full matching [8, 48, 49, 47]. These algorithms often come with guaranteed bounds on optimality, frequently incorporating geometric consistency constraints.
Partial-to-Full Shape Matching.
[46] demonstrate that for partial-to-full matching, the functional map has a slanted diagonal structure, which has, for example, been used in [37] to solve multi-part partial-to-partial correspondence; however, under the assumption that a full template is given. Several methods also operate in the spectral domain [3, 14, 15] in the learned setting or without the functional map layer [11, 12] and explicitly tackle the partial-to-full case.
Partial-to-Partial Shape Matching.
Few works thus far explicitly address the partial-to-partial matching setting, which is particularly challenging as the overlapping region between shapes is unknown. As such, the overlapping region must simultaneously be identified and matched under an unknown deformation. DPFM [3] demonstrates that spectral methods are applicable in this domain using an additional attention mechanism, albeit using ground-truth supervision for feature learning. In SM-COMB, Roetzer et al. [48] formulated shape matching as an integer linear program (ILP) over the space of orientation-preserving diffeomorphisms, demonstrating that the latter constitute strong priors to guide partial-to-partial matching. However, holes must be synthetically filled to obtain a closed manifold. Recently, Ehm et al. [24] formulated a geometrically consistent partial-to-partial shape matching nonlinear integer program named GC-PPSM. While both SM-COMB and GC-PPSM are compatible with various surface descriptors, they achieve SOTA results only with high-quality pre-trained input features (e.g. from URSSM [15]), which can be prohibitive when these are not available  [48, 24].
2.3 Feature Descriptors in Shape Correspondence
Many axiomatic and learning-based methods for partial shape matching rely on surface feature descriptors as guidance for correspondence computation. Feature descriptors, like SHOT [53], can be handcrafted based on extrinsic quantities like XYZ-coordinates or surface normals. Many surface descriptors are sensitive to surface degradation, such as holes, or are not invariant under rotations, as is the case for XYZ coordinates. The common intrinsic heat kernel (HKS) [52] and wave kernel signatures (WKS) [5] are invariant under isometric deformations; however, they are not effective under partiality as the isometry assumption can be heavily violated. On the other hand, methods like DiffusionNet [51] aim to learn robust surface descriptors from data while still using some features like XYZ coordinates or HKS descriptors as input. Considering that 3D shapes of many datasets are typically oriented in an upright manner [10, 2, 13, 45, 22], or generally positioned in a canonical pose [26] this can lead to biased predictions. Moreover, learned features require large training datasets, and a high-quality (self-)supervision signal cannot be guaranteed under severe partiality.
Figure 2: Left: We use a collection of multiple complete geometry datasets to generate novel partial shape pairs. Correspondences are propagated based on our manually created cross-category and cross-dataset correspondences (black lines, see Section 3.1) and existing intra-dataset correspondences (blue dashed lines). Right: Given an input shape with full geometry from our shape collection, we first re-mesh the shape, then apply ray casting to simulate a unidirectional 3D scanning procedure to introduce partiality (see Section 3.2), and finally apply a deterministic random rotation.
3 Cross-Category and Cross-Dataset Procedural Data Generation
To overcome the limitations in realism and sample size of existing partial shape benchmarks, we propose a novel framework enabling the procedural generation of partial shape instances with ground-truth correspondences based on a given collection of 3D shapes. Our framework supports shapes from multiple categories to generate non-isometric partial shape pairs. Further, we propose linking correspondences between multiple datasets to vastly increase the diversity and representativeness. We use ray casting to generate more realistic partial shapes than benchmarks based on synthetic planar cuts and holes. Furthermore, our pipeline supports re-meshing while preserving shape correspondences within a collection and thus foster discretisation-agnostic method development. We summarise the main idea in Fig. 2.
Our pipeline is customisable and versatile in generating novel full-to-full, partial-to-full and partial-to-partial correspondence pairs. In the main body of this paper, we focus on the under-explored and in our view most novel partial-to-partial and partial-to-full settings. A detailed description of our data generation in the supp. mat. Sec. 10. All code will be publicly available.
Refer to caption Refer to caption Refer to caption Refer to caption
Source Targets
Figure 3: For the first time we define missing parts in full shapes. As the horse does not have a human upper body or horns, they are defined as missing parts in the deer, goat and centaur.
3.1 Cross-Dataset and Cross-Category Correspondences
Current datasets commonly lack correspondences between different categories, limiting data availability and diversity [13, 36]. We alleviate this shortcoming by manually establishing correspondences across categories and across different datasets. Therefore, we treat each category that does not have cross-category correspondences as a separate entity. Then, we select a template shape for each entity and manually choose a set of corresponding landmarks on pairs of templates using FaceForm [28] (see Sec. 10). We then morph one template into a second to establish dense point-to-surface correspondences (in barycentric coordinates). Using the tightly aligned shapes, we obtain dense correspondences by projecting points onto triangles of the other surface. Afterwards, we propagate the dense correspondences to the entire shape category using the existing intra-category correspondences. Overall, we build a spanning tree of pairwise dense correspondences that covers all possible shape templates (cross-category and cross-dataset). A detailed tree is presented on our project page (see link in abstract).
3.2 Partiality Generation
Since dense correspondences can be generated for all possible shape pairs as described in Section 3.1, random pairs can be selected to generate new examples with varying degrees of partiality. We require the shapes to have a semantic overlap; for example, a human shape can not be matched to an elephant. To generate a partial-to-partial shape matching pair, we first sample two corresponding shapes (denoted as $\mathcal{X}$ and $\mathcal{Y}$) from the aggregated dataset. For a partial-to-full shape matching pair, shape $\mathcal{X}$ is paired with a template shape from the same category.
Partiality via Ray Casting.
Given a shape $\mathcal{X}$, we introduce partiality by simulating a simple unidirectional 3D scanning procedure based on ray casting. Doing so leads to a variety of different boundary shapes and overlapping behaviour (see Fig. 1 for examples). To perform ray casting, we first centre and scale each full shape $\mathcal{X}$ to be within a fixed unit bounding box and then choose random camera poses on the unit sphere. Partiality is obtained by casting a bundle of $R\times R$ rays from the sampled camera pose onto the full shape $\mathcal{X}$, where $R$ is the number of rays along the x-axis. All hit triangles forming the largest connected component are kept to be part of the partial shape $\hat{\mathcal{X}}$. Finally, each partial shape $\hat{\mathcal{X}}$ is scaled and translated back to its original size and location. Similarly, we get partial shape $\hat{\mathcal{Y}}$ from $\mathcal{Y}$.
Random camera poses could lead to non-existent or prohibitively small overlap in the partial-to-partial setting. Therefore, we sample two camera poses with the constraint that their maximum angular disparity in both azimuth and elevation does not exceed $\alpha$ degrees. This is to generate an overlap of a certain size to ensure meaningful matching instances. For the overlap computation, we rigidly align the shapes $\mathcal{X}$ and $\mathcal{Y}$ using the given correspondences to ensure that both partial shapes are semantically similar and perform the previously described partiality generation on both shapes. Subsequently, we quantify the overlap as the percentage of vertices from $\hat{\mathcal{X}}$ that are mapped to $\hat{\mathcal{Y}}$ and vice-versa. If both overlaps do not fall within a predefined range, the process iterates with new random camera poses until the overlap meets the specified criteria, or until a maximum of $m$ iterations is reached. In the latter case, the partial shapes whose overlap is the closest to the desired range are selected (see Sec. 10 in supp. mat. for more details.
Semantic Partiality.
Since we combine multiple categories and datasets, some pairs of shapes cannot be semantically matched (e.g, a deer’s horn can be matched to a horse). To this end, we define for the first time missing parts in shapes as seen in Fig. 3. This creates a novel objective for shape matching methods that warrants future research.
3.3 Virtually Infinite Generation of Shape Matching Instances
Existing partial shape datasets provide a precomputed and fixed number of holes/cuts, creating partial shape pairs from identical meshes in varying poses. Instead, our proposed framework can generate partial shapes from a larger and more diverse pool of shapes, allowing for pairing any compatible shapes from different categories, potentially originating from different datasets. This creates a more challenging scenario encompassing simultaneous non-isometric deformations and partiality, representing an important step towards more realistic settings. Furthermore, our framework introduces more variety by using different camera poses for ray casting and re-meshing shapes to different resolutions.
4 The BeCoS Benchmark
In addition to the procedural data generation framework, we propose the BeCoS benchmark, which is a specific instantiation of our general framework introduced in Sec. 3. BeCoS contains realistic partial shapes of established datasets with cross-category and cross-dataset correspondences. The purpose of BeCoS is to challenge current state-of-the-art partial shape matching methods and stimulate progress in the field through a standardized evaluation. To create our benchmark, we apply the pipeline introduced in Sec. 3 on a combination of several existing datasets.
Datasets.
We include seven different datasets with four-legged creatures and humanoid shapes encompassing challenging and diverse shape matching scenarios, which exhibit both isometric and non-isometric deformations: TOSCA [13], FAUST [10], SCAPE [2], KIDS [45], DT4D [39], SMAL [56] and SHREC’20 [22]. From these datasets, we use a total of 2543 shapes to generate pairwise matching problems. We partition our dataset into predefined training, validation, and test splits, ensuring evaluation is performed on unseen categories. To limit the imbalance in dataset size, we oversample the smaller datasets in our train set The number of shape matching instances in the train/test/validation splits are 10185/142/137. We limit the test set size to easily compare to axiomatic methods, which are often computationally heavy.
Predefined dataset parameters.
We choose the following parameters in the data generation pipeline to create a challenging but reasonable benchmark. For re-meshing, we fix the range of vertices to a random number between $9000{-}10{,}000$ to facilitate training of current supervised/unsupervised methods and limit memory consumption. Our main setting introduces a random rotation around a single axis, assuming each shape has a natural ‘upright’ direction. For the partial-to-partial setting we set the maximum camera angle to $\alpha=\frac{\pi}{4}$. Similar to [3] and [24], we aim to have an overlap of $10\%-90\%$, with maximum iterations $m$ set to 10 (see supp. mat. Sec. 8.2 for distribution of overlap).
We aim to generate a challenging but feasible configuration, which provokes the shortcomings of state-of-the-art methods while still being achievable. In the future, once the benchmark is saturated due to the development of accurate partial shape matching methods, our customisable framework enables a straight-forward generation of more challenging variants.
Methods Type F2F P2F P2P Input
features
GeomFMaps [21] Sup wks, xyz
DPFM [3] Sup/Unsup wks, xyz
ULRSSM [15] Unsup wks, xyz
Smooth Shells [25] Axio shot
PFM [46] Axio shot
SM-COMB [48] Axio xyz
GC-PPSM  [24] Axio xyz
Table 2: Overview of baselines with their methodology type (Supervised,Unsupervised or Axiomatic), their possible problem statements (P2P: Partial-to-Partial, P2F: Partial-to-Full and F2F:Full-to-Full) and their corresponding input features. All default values are used for all baselines to produce our results.
Full-to-Full (F2F) Source GT GeomFMaps ULRSSM Smooth Shells

### v2

@@ Abstract
Finding correspondences between 3D deformable shapes is an important and long-standing problem in geometry processing, computer vision, graphics, and beyond. While various shape matching datasets exist, they are mostly static or limited in size, restricting their adaptation to different problem settings, including both full and partial shape matching. In particular the existing partial shape matching datasets are small (fewer than 100 shapes) and thus unsuitable for data-hungry machine learning approaches. Moreover, the type of partiality present in existing datasets is often artificial and far from realistic. To address these limitations, we introduce a generic and flexible framework for the procedural generation of challenging full and partial shape matching datasets. Our framework allows the propagation of custom annotations across shapes, making it useful for various applications. By utilising our framework and manually creating cross-dataset correspondences between seven existing (complete geometry) shape matching datasets, we propose a new large benchmark BeCoS with a total of 2543 shapes. Based on this, we offer several challenging benchmark settings, covering both full and partial matching, for which we evaluate respective state-of-the-art methods as baselines. Visualisations and code of our benchmark can be found at: https://nafieamrani.github.io/BeCoS/.
@@ 1 Introduction
Finding correspondences between deformable shapes is a fundamental problem in geometry processing, with various applications such as texture and deformation transfer [dinh2005texture, ovsjanikov2012functional], or statistical shape analysis [hasler2009statistical, bogo2014faust, elamrani2024fuss]. The problem has been widely studied in the literature  [ovsjanikov2012functional, eisenberger2020, abdelreheem2023zero, wimmer2024back, cao2023unsupervised], and can vary significantly in complexity. For example, current methods still struggle when estimating severe non-isometric deformations for shapes of different categories [dyke2020shrec, panine2022non, eisenberger2020, bastian2023hybrid, cao2024spectral], or when they are only partially observed [roetzer2022scalable, ehm2024partial]. Current non-isometric shape datasets are often limited in size and flexibility [Zuffi:CVPR:2017, melzi2019shrec, dyke2020shrec, magnet2022smooth, li20214dcomplete], restricting their ability to support dense correspondence propagation or custom annotations, which reduces their applicability to broader tasks.
@@
In real-world scenarios, it is particularly common to encounter shapes with missing parts due to occlusions, limited sensor coverage, or acquisition errors during 3D scanning. The setting in which at least part of one shape does not exist on the other is called partial shape matching [rodola2017partial, attaiki2021dpfm, roetzer2022scalable, ehm2023geometrically, ehm2024partial]. Despite its practical relevance, partial shape matching has received limited attention in the context of machine learning-based approaches, partly because of the increased difficulty, but largely due to the lack of suitable and realistic large-scale datasets, a requirement for data-hungry learning algorithms. The full-to-full matching setting is not only simpler but also benefits from wider data availability which has supported the development of data-driven methods that already display convincing performance, even without supervision [halimi2019unsupervised, roufosse2019unsupervised, cao2023unsupervised]. In contrast, existing datasets for partial shape correspondence often suffer from a small number of shapes, limited realism of generated partiality (e.g., cuts or holes) [cosmo2016shrec], and a lack of diversity in shape categories [bracha2023partial, attaiki2021dpfm]. This limited data availability hampers the progress of data-driven methods for partial shape matching; recent approaches have thus required pre-training with large full-geometry datasets [cao2023unsupervised], or rely on ground-truth correspondences [attaiki2021dpfm] which are challenging to obtain in practice.
@@
To address the absence of large datasets and foster research on (partial) shape correspondence, we introduce a procedural data generation framework to generate diverse shape pairs from a shape network. By simulating more realistic partiality patterns using ray casting, we bridge the gap between existing synthetic datasets and real-world partial shape correspondence problems. Furthermore, we propose BeCoS, a carefully curated benchmark based on existing datasets comprising diverse shape categories, thus enabling to assess the performance of existing and foster the development of novel methods. Through extensive experiments of current state-of-the-art partial and full shape matching methods, we demonstrate the limitations of partial methods and highlight the need for further research in this area.
@@
  • Shape Network for Correspondence Propagation: We unify multiple existing mesh-based datasets through manual cross-category annotations, constructing a large and diverse shape network that indicates between which shapes correspondences exist. This shape network facilitates the propagation of dense correspondences and custom annotations across shapes, enhancing its applicability to various shape analysis tasks.
  • Scalable Framework for Partial Shape Correspondence: We introduce a flexible and scalable framework that enables the generation of training examples for partial shape correspondence tasks that is applicable to new datasets.
  • Benchmark and Performance Analysis: As a concrete realization of our framework, we present a challenging non-rigid shape dataset, serving as a standardised benchmark for shape matching methods. We further conduct an extensive evaluation of state-of-the-art techniques on this benchmark, offering insights into their strengths, limitations, and key challenges. Our results indicate that partial shape matching remains an open problem, motivating future research in the field.
@@ 2.1 Datasets
Although several 3D shape datasets exist, we align with established deformable shape matching literature by focusing on humanoid and four-legged animal shapes. These categories support continuous deformations and offer well-defined dense correspondences, making them particularly suitable for our study. In the following, we provide an overview of existing datasets for full-to-full (F2F), partial-to-full (P2F), and partial-to-partial (P2P) shape matching datasets. Table 1 summarises existing deformable shape matching datasets.
@@
Full-to-Full Datasets. Full-to-full shape correspondence datasets, such as FAUST [bogo2014faust] and SCAPE [anguelov2005scape], have been essential in methodological progress on shape correspondence. They encompass humans in various poses, initially emphasising point-wise correspondence recovery under isometric deformations. SHREC’19 [melzi2019shrec] was introduced to evaluate the generalisation capabilities of methods to different mesh connectivity on human scans. Datasets like TOSCA [bronstein2008numerical], SMAL [Zuffi:CVPR:2017], SHREC’20 [dyke2020shrec], and DeformingThings4D [magnet2022smooth, li20214dcomplete] have shifted the emphasis towards non-isometric correspondence between various animal categories and humanoid figurines. Although deep learning methods have achieved state-of-the-art performance in shape correspondence, non-isometric datasets remain relatively small and static, limiting their applicability in training and evaluating deep learning-based correspondence methods. While full-to-full shape matching has received considerable attention, the partial setting has been explored less (see Table 1).
Setting Dataset # Unique Shape Non
Shapes Type Isometry